# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

### v1

@@ 1 Introduction and Main Results
Furthermore, there exists a constant $C=C(G,T,\mathcal{A},\mathcal{A}_{\Gamma})>0$ such that
$\begin{aligned} &\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))}+\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})}+\|(Z,\widehat{Z})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2})}\\ &\leq C\left(\|(z_{T},z_{\Gamma,T})\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})}+\|F_{1}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}+\|F\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G;\mathbb{R}^{N}))}\right.\\ &\hskip 28.45274pt\left.+\|F_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}+\|F_{\Gamma}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma;\mathbb{R}^{N}))}\right).\end{aligned}$
We recall the following well-known technical result due to Fursikov-Imanuvilov (see [15]).
@@
Since $|\nabla\psi|^{2}=|\nabla_{\Gamma}\psi|^{2}+|\partial_{\nu}\psi|^{2}$ on $\Gamma$, the function $\psi$ in Lemma 1.1 satisfies
$\nabla_{\Gamma}\psi=0,\qquad|\nabla\psi|=|\partial_{\nu}\psi|,\qquad\partial_{\nu}\psi\leq-c<0\quad\text{on}\;\,\Gamma,\quad\text{for some constant }\,c>0.$ (1.3)
For any parameters $\lambda,\mu\geq 1$, we choose the following weight functions:
$\alpha\equiv\alpha(t,x)=\frac{e^{\mu\psi(x)}-e^{2\mu|\psi|_{\infty}}}{t(T-t)},\quad\ell=\lambda\alpha,\quad\theta=e^{\ell},\quad\varphi\equiv\varphi(t,x)=\frac{e^{\mu\psi(x)}}{t(T-t)},$ (1.4)
for any $(t,x)\in(0,T)\times\overline{G}$. Moreover, the functions $\alpha$ and $\varphi$ are constant on the boundary $\Gamma$, so that
$\nabla_{\Gamma}\alpha=\nabla_{\Gamma}\varphi=0\quad\text{on}\;\,\Gamma.$
It is easy to verify that for some suitable constant $C=C(G)>0$, the following inequalities hold:
$\begin{aligned} &\,\varphi\geq CT^{-2},\qquad|\varphi_{t}|\leq CT\varphi^{2},\qquad|\varphi_{tt}|\leq CT^{2}\varphi^{3},\\ &|\alpha_{t}|\leq CTe^{2\mu|\psi|_{\infty}}\varphi^{2},\qquad|\alpha_{tt}|\leq CT^{2}e^{2\mu|\psi|_{\infty}}\varphi^{3},\end{aligned}$ (1.5)
for all $(t,x)\in(0,T)\times\overline{G}$. In the following, we choose the set $G_{1}$ from Lemma 1.1 such that $G_{1}\Subset G_{0}$.
The first main result of this paper is the following Carleman estimate for the solutions of (1.2).
@@
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}\,dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\,d\sigma dt\\ &+\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}\,dxdt+\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}\,dxdt+\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+\mu\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}\,dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|F_{\Gamma}|^{2}\,d\sigma dt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}\,dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}\,d\sigma dt\Bigg],\end{aligned}$ (1.6)
for all $\mu\geq\mu_{0}$ and $\lambda\geq\lambda_{0}(e^{2\mu|\psi|_{\infty}}T+T^{2})$.
In [6, Theorem 4.1], a Carleman estimate for equation (1.2) is presented in the absence of divergence source terms (i.e., $F\equiv F_{\Gamma}\equiv 0$), where only Laplacian operators are involved (i.e., $\mathcal{A}\equiv\mathcal{A}_{\Gamma}\equiv I_{N}$), and where the parameters $\lambda$ and $\mu$ do not depend on the final control time $T$. This estimate is derived using the weighted identity method, applying two weighted identities for the stochastic parabolic operators $\text{``}dz+\Delta z\,dt\text{''}$ and $\text{``}dz_{\Gamma}+\Delta_{\Gamma}z_{\Gamma}\,dt-\partial_{\nu}z\,dt\text{''}$. In this paper, we extend this result to address the presence of divergence source terms and general second-order parabolic operators. Our goal is to derive the new Carleman estimate, given in Theorem 1.1. To obtain this estimate, we combine two key methods. First, we apply the weighted identity method to the operators $\text{``}dz+\nabla\cdot(\mathcal{A}\nabla z)\,dt\text{''}$ and $\text{``}dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial^{\mathcal{A}}_{\nu}z\,dt\text{''}$, refining the known Carleman estimate from [6, Theorem 4.1] to a more general form. This generalization includes diffusion terms represented by the matrices $\mathcal{A}$ and $\mathcal{A}_{\Gamma}$, and explicitly incorporates the dependence of $\lambda$ and $\mu$ on the final control time $T$. Second, to handle the divergence source terms $F$ and $F_{\Gamma}$, we employ the duality technique introduced in [23] for the stochastic setting, which ultimately leads to the desired Carleman estimate (1.6). For more details on the derivation of Carleman estimates for stochastic parabolic equations with different types of boundary conditions, we refer the reader to [3, 14, 23, 32] and [27, Chapter 9] for the case of Dirichlet boundary conditions, to [7, 34] for the case of Neumann and Robin boundary conditions, and to [4, 5, 6] for the case of dynamic boundary conditions. For results on Carleman estimates for some coupled forward-backward stochastic parabolic equations, see [23, 29].
As an application of our Carleman estimate (1.6), we study the null and approximate controllability problems for the following forward stochastic parabolic equation
@@
$\begin{cases}\begin{array}[]{ll}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=(a_{1}y+B_{1}\cdot\nabla y+\mathbbm{1}_{G_{0}}u)\,dt+v_{1}\,dW(t)&\text{in}\,\,Q,\\ dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma})\,dt+v_{2}\,dW(t)&\text{on}\,\,\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\text{on}\,\,\Sigma,\\ (y,y_{\Gamma})|_{t=0}=(y_{0},y_{\Gamma,0})&\text{in}\,\,G\times\Gamma,\end{array}\end{cases}$ (1.7)
where $(y,y_{\Gamma})$ is the state variable, $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$ is the initial state, the coefficients $a_{1}$, $a_{2}$, $B_{1}$, and $B_{2}$ are bounded and adapted stochastic processes, i.e.,
$a_{1}$ $\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(G)),\quad B_{1}\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(G;\mathbb{R}^{N})),$
$a_{2}$ $\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(\Gamma)),\quad B_{2}\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(\Gamma;\mathbb{R}^{N})).$
In (1.7), the control function consists of the triple
@@
From [4, Theorem 2.1], we deduce that (1.7) is well-posed, i.e., for any $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$ and $(u,v_{1},v_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, the system (1.7) admits a unique weak solution
$(y,y_{\Gamma})\in L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1}),$
such that for some positive constant $C=C(G,T,\mathcal{A},\mathcal{A}_{\Gamma})$, we have
$\,\|(y,y_{\Gamma})\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))}+\|(y,y_{\Gamma})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})}$
$\leq C\Big(\|(y_{0},y_{\Gamma,0})\|_{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})}+\|u\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))}$
$\hskip 28.45274pt+\|v_{1}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}+\|v_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}\Big).$

### v2

@@ 1 Introduction and Main Results
Moreover, there exists a constant $C>0$, depending only on $G$, $T$, $M$, and $M_{\Gamma}$, such that
$\max_{0\leq t\leq T}\mathbb{E}\|(z(t),z_{\Gamma}(t))\|_{\mathbb{L}^{2}}^{2}+\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})}^{2}+\|(Z,\widehat{Z})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2})}^{2}$
$\quad\leq C\bigg(\mathbb{E}\|(z_{T},z_{\Gamma,T})\|_{\mathbb{L}^{2}}^{2}+\|F_{1}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}^{2}+\sum_{j=1}^{N}\|F_{j}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}^{2}$
$\qquad\qquad+\|F_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}^{2}+\sum_{j=1}^{N}\|{F_{\Gamma}}_{j}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}^{2}\bigg).$
In order to establish our main Carleman estimate for equation (1.2), we first recall the following classical technical result due to Fursikov and Imanuvilov (see [19]).
@@
Since $|\nabla\psi|^{2}=|\nabla_{\Gamma}\psi|^{2}+|\partial_{\nu}\psi|^{2}$ on $\Gamma$, the function $\psi$ in Lemma 1.1 also satisfies
$\nabla_{\Gamma}\psi=0,\qquad|\nabla\psi|=|\partial_{\nu}\psi|,\qquad\partial_{\nu}\psi\leq-C<0\quad\textnormal{on}\;\,\Gamma,\quad\textnormal{for some constant }\,C>0.$ (1.5)
For any parameters $\lambda,\mu\geq 1$, we introduce the following weight functions:
$\begin{aligned} \gamma\equiv\gamma(t)&:=(t(T-t))^{-1},\qquad\alpha\equiv\alpha(t,x):=\bigl(e^{\mu\psi(x)}-e^{2\mu\|\psi\|_{\infty}}\bigr)\gamma(t),\\ \ell&:=\lambda\alpha,\qquad\theta:=e^{\ell},\qquad\varphi\equiv\varphi(t,x):=e^{\mu\psi(x)}\,\gamma(t),\end{aligned}$ (1.6)
for all $(t,x)\in(0,T)\times\overline{G}$. Note that $\alpha$ and $\varphi$ are constant on the boundary $\Gamma$, which implies
$\nabla_{\Gamma}\alpha=\nabla_{\Gamma}\varphi=0\quad\textnormal{on}\;\,\Gamma.$
Moreover, one can readily verify that there exists a constant $C=C(G)>0$ such that
$\begin{aligned} &\varphi\geq 4T^{-2},\qquad|\varphi_{t}|\leq CT\varphi^{2},\qquad|\varphi_{tt}|\leq CT^{2}\varphi^{3},\\ &|\alpha_{t}|\leq CTe^{2\mu\|\psi\|_{\infty}}\varphi^{2},\qquad|\alpha_{tt}|\leq CT^{2}e^{2\mu\|\psi\|_{\infty}}\varphi^{3}.\end{aligned}$ (1.7)
In what follows, we choose the weight functions $\theta$, $\ell$, and $\varphi$ as in (1.6), and the function $\psi$ as in Lemma 1.1, with $G_{1}$ being any fixed nonempty open subset such that $G_{1}\Subset\mathcal{B}$, for any nonempty open subset $\mathcal{B}\Subset G$. The first main result of this paper is the following fundamental Carleman estimate for equation (1.2), stated as follows:
@@
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}|z|^{2}\,dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}|z_{\Gamma}|^{2}\,d\sigma dt\\ &+\lambda\mu\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}\,dxdt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma dt\\ &\leq C\bigg[\lambda^{3}\mu^{4}\mathbb{E}\int_{0}^{T}\int_{\mathcal{B}}\theta^{2}\varphi^{3}|z|^{2}\,dxdt+\mathbb{E}\iint_{Q}\theta^{2}|F_{1}|^{2}\,dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}|F_{2}|^{2}\,d\sigma dt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\sum_{j=1}^{N}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F_{j}|^{2}\,dxdt+\lambda^{2}\mu^{2}\sum_{j=1}^{N}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|{F_{\Gamma}}_{j}|^{2}\,d\sigma dt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|Z|^{2}\,dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|\widehat{Z}|^{2}\,d\sigma dt\bigg],\end{aligned}$ (1.8)
for all $\mu\geq\mu_{0}$ and $\lambda\geq\lambda_{0}(e^{2\mu\|\psi\|_{\infty}}T+T^{2})$.
To the best of our knowledge, this paper is the first to establish the Carleman estimate (1.8) for backward anisotropic stochastic parabolic equations with $\mathbb{L}^{2}$ bulk–surface source terms and $\mathbb{H}^{-1}$ bulk–surface source terms appearing as weak divergence-type sources, coupled with the general dynamic boundary condition
$\begin{cases}dz_{\Gamma}+\displaystyle\sum_{j,k=1}^{N}D_{j}\Big(a_{\Gamma}^{jk}(x)D_{k}z_{\Gamma}\Big)\,dt-\displaystyle\sum_{j,k=1}^{N}a^{jk}(x)\frac{\partial z}{\partial x_{j}}\nu^{k}\,dt=\Big(F_{2}-\displaystyle\sum_{j=1}^{N}F_{j}\nu^{j}+\sum_{j=1}^{N}D_{j}{F_{\Gamma}}_{j}\Big)\,dt+\widehat{Z}\,dW(t),&\textnormal{on }\Sigma,\\[2.84526pt] z_{\Gamma}=z|_{\Gamma},&\textnormal{on }\Sigma.\end{cases}$
Such systems are substantially more challenging to analyze than classical stochastic parabolic equations with standard Dirichlet boundary conditions (see [4, 29, 40] and [34, Chapter 9]), as well as Neumann or Robin boundary conditions (see [10, 43]). The additional difficulties arise from the combined presence of bulk–surface interactions, weak divergence-type source terms, and dynamic boundary effects, which induce strong coupling mechanisms and reduced regularity.
In the existing literature, Carleman estimates for systems with dynamic boundary conditions are available only in limited and simplified settings (see [5, 6, 7]) or for certain coupled forward–backward stochastic systems (see [3, 38]). However, a comprehensive Carleman estimate for backward anisotropic stochastic parabolic equations with general $\mathbb{H}^{-1}$ bulk–surface couplings and dynamic boundary conditions, together with an explicit Carleman constant independent of the final time $T$, has not yet been established. Addressing this gap is the main objective of the present paper.
In addition, we apply the derived Carleman estimate (1.8) to study the null and approximate controllability, as well as an insensitizing control problem, for forward reaction–convection–diffusion stochastic systems with general dynamic boundary conditions. The explicit dependence of our Carleman constant on $T$ is exploited, for the first time, to obtain a sharp estimate on the problem data for the cost of null controllability of the associated controlled adjoint forward stochastic problem. In particular, this allows us to estimate the minimal norm of the controls required to drive the system to zero at time $T$. These results highlight both the novelty and the significance of the present work.
By establishing the Carleman estimate (1.8), we address three fundamental challenges:
  1. 1.
    We consider a fully general class of stochastic parabolic operators involving anisotropic diffusion matrices $(a^{jk})_{1\leq j,k\leq N}$ in the bulk and $(a_{\Gamma}^{jk})_{1\leq j,k\leq N}$ on the boundary, thus extending beyond the isotropic diffusion framework commonly studied in the literature. This generality introduces many additional terms that must be treated with care.
  2. 2.
    We derive an explicit dependence of the Carleman constant on the final time $T$, which is essential for quantitative estimates and plays a key role in controllability problems.
  3. 3.
    We allow source terms in the drift components to belong to negative Sobolev spaces, which significantly relaxes the regularity assumptions and enables the treatment of weak divergence-type sources both in the bulk and on the boundary.
To accomplish these objectives, we combine two complementary analytical methods.
  • First, we apply the weighted identity method to the following general stochastic parabolic-type operators:
    $\textnormal{``}dz+\sum_{j,k=1}^{N}\frac{\partial}{\partial x_{j}}\Big(a^{jk}(x)\frac{\partial z}{\partial x_{k}}\Big)\,dt\textnormal{''}\quad\textnormal{and}\quad\textnormal{``}dz_{\Gamma}+\sum_{j,k=1}^{N}D_{j}\Big(a_{\Gamma}^{jk}(x)D_{k}z_{\Gamma}\Big)\,dt-\sum_{j,k=1}^{N}a^{jk}(x)\frac{\partial z}{\partial x_{j}}\nu^{k}\,dt\textnormal{''}.$
    As a result, our Carleman estimate extends Theorem 4.1 of [7], which corresponds to the particular case
    $(a^{jk})_{N\times N}=(a_{\Gamma}^{jk})_{N\times N}=(\delta^{jk})_{N\times N}=I_{N},\qquad F_{j}={F_{\Gamma}}_{j}=0,\quad j=1,2,\cdots,N,$
    where $I_{N}$ denotes the $N\times N$ identity matrix and $\delta^{jk}$ is the Kronecker delta. In addition, we obtain an explicit dependence of the Carleman constant on the final time $T$. This generalization allows for anisotropic diffusion effects through the matrices $(a^{jk})_{N\times N}$ and $(a_{\Gamma}^{jk})_{N\times N}$, thus covering a substantially broader class of diffusion processes. A major difficulty in our analysis arises from the treatment of new boundary integral terms generated by the general dynamic boundary condition
    $\begin{cases}dz_{\Gamma}+\displaystyle\sum_{j,k=1}^{N}D_{j}\Big(a_{\Gamma}^{jk}(x)D_{k}z_{\Gamma}\Big)\,dt-\displaystyle\sum_{j,k=1}^{N}a^{jk}(x)\frac{\partial z}{\partial x_{j}}\nu^{k}\,dt=F_{2}\,dt+\widehat{Z}\,dW(t),&\textnormal{on }\Sigma,\\[2.84526pt] z_{\Gamma}=z|_{\Gamma},&\textnormal{on }\Sigma.\end{cases}$
    These terms require a careful and adapted analysis, especially in the proof of Lemma 3.3.
  • Second, to handle the weak divergence-type source terms $F_{j}$ and ${F_{\Gamma}}_{j}$, $j=1,2,\cdots,N$, we employ a duality method combined with suitable optimization arguments (see [29] and the references therein for the case of Dirichlet boundary conditions). This strategy allows us to absorb the divergence-type contributions into the Carleman inequality and thus derive the desired estimate (1.8). In contrast to [29], where Carleman estimates are mainly used to relax regularity assumptions on diffusion source terms, our approach is specifically designed to address weak divergence-type sources of the form
    $\sum_{j=1}^{N}\frac{\partial F_{j}}{\partial x_{j}}\quad\textnormal{in }Q,\qquad\textnormal{and}\qquad\sum_{j=1}^{N}D_{j}{F_{\Gamma}}_{j}\quad\textnormal{on }\Sigma.$
    The main difficulty lies in the fact that the functions $F_{j}$ and ${F_{\Gamma}}_{j}$ belong only to $L^{2}$-spaces. Consequently, the corresponding divergence terms are only defined in the distributional sense and belong to the negative Sobolev space $H^{-1}$, which prevents the direct application of standard Carleman techniques and necessitates the duality-based framework developed in this work.
In this paper, as an application of our Carleman estimate (1.8), we investigate two control problems: the null controllability and an insensitizing control problem for the following forward anisotropic stochastic reaction–convection–diffusion system with general dynamic boundary conditions:
@@
$\begin{cases}\begin{array}[]{ll}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=\bigl(\xi+a_{1}y+B_{1}\cdot\nabla y+\mathbbm{1}_{G_{0}}u\bigr)\,dt+v_{1}\,dW(t)&\textnormal{in }Q,\\[3.00003pt] dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=\bigl(\xi_{\Gamma}+a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma}\bigr)\,dt+v_{2}\,dW(t)&\textnormal{on }\Sigma,\\[3.00003pt] y_{\Gamma}=y|_{\Gamma}&\textnormal{on }\Sigma,\\[3.00003pt] (y,y_{\Gamma})|_{t=0}=(y_{0}+\tau_{1}\widehat{y}_{0},y_{\Gamma,0}+\tau_{2}\widehat{y}_{\Gamma,0})&\textnormal{in }G\times\Gamma.\end{array}\end{cases}$ (1.9)
Here, $(y,y_{\Gamma})$ denotes the state variable of the system, with $y$ representing the bulk state and $y_{\Gamma}$ its trace on the boundary. The initial datum $(y_{0},y_{\Gamma,0})$ is assumed to belong to $L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$. We further assume that $(\widehat{y}_{0},\widehat{y}_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$ is an unknown perturbation of $(y_{0},y_{\Gamma,0})$ satisfying
$\mathbb{E}\,\|(\widehat{y}_{0},\widehat{y}_{\Gamma,0})\|_{\mathbb{L}^{2}}^{2}=1,$
and that $\tau_{1},\tau_{2}\in\mathbb{R}$ are unknown small parameters. The functions $\xi\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$ and $\xi_{\Gamma}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$ denote the given source terms in the bulk and on the boundary, respectively. Accordingly, the term $(\tau_{1}\widehat{y}_{0},\tau_{2}\widehat{y}_{\Gamma,0})$ may be regarded as a small perturbation of the nominal initial data $(y_{0},y_{\Gamma,0})$. The coefficients $a_{1}$ and $B_{1}$ describe the reaction and convection effects in the bulk, while $a_{2}$ and $B_{2}$ account for the corresponding effects on the boundary. They are assumed to be bounded and $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted stochastic processes. More precisely, we assume that
$a_{1}$ $\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(G)),\qquad\quad B_{1}\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(G;\mathbb{R}^{N})),$
$a_{2}$ $\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(\Gamma)),\qquad\quad B_{2}\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(\Gamma;\mathbb{R}^{N})).$
The control mechanism in (1.9) consists of a distributed control $u$ acting on the interior subdomain $G_{0}\Subset G$, together with stochastic controls $v_{1}$ and $v_{2}$ acting in the bulk and on the boundary, respectively. Accordingly, the admissible control triple satisfies
@@
Similarly to the proof of [5, Theorem 2.1], one can show that system (1.9) is well posed. More precisely, for any initial datum $(y(0),y_{\Gamma}(0))\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$, source terms $\xi\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$ and $\xi_{\Gamma}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, and any control triple
$(u,v_{1},v_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)),$
system (1.9) admits a unique weak solution
$(y,y_{\Gamma})\in L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1}).$
Moreover, there exists a constant $C=C(G,T,\beta_{0},M,M_{\Gamma})>0$ such that
$\max_{0\leq t\leq T}\mathbb{E}\,\|(y(t),y_{\Gamma}(t))\|^{2}_{\mathbb{L}^{2}}+\|(y,y_{\Gamma})\|^{2}_{L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})}$
$\qquad\leq C\Big(\mathbb{E}\,\|(y(0),y_{\Gamma}(0))\|^{2}_{\mathbb{L}^{2}}+\|\xi\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}+\|\xi_{\Gamma}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}$
[…]